Is (9, -1) a solution of x + 5y = 4?

Answer by nerdybill(7384)

Is (9, -1) a solution of x + 5y = 4?
since they give you an (x,y) pair, this defines:
x as 9
and
y as -1
plug it into:
x + 5y = 4
9 + 5(-1) = 4
9 + (-5) = 4
9 - 5 = 4
4 = 4 (TRUE)
Since, the statment is true, we conclude that:
(9, -1) IS a solution